Determine the pH of each of the following solutions.



(a) 0.213 M boric acid (weak acid with Ka = 5.8e-10).





(b) 0.394 M phenol (weak acid with Ka = 1.3e-10).





(c) 0.869 M pyridine (weak base with Kb = 1.7e-09).
Answer #1

Having the general reaction:

Acid: HA + H2O = A- + H3O +

Base: A- + H2O = HA + OH-

You have Ka's expression for:

a) Ka = [A-] * [H3O +] / [HA]

5.8x10 ^ -10 = X ^ 2 / 0.213

It clears X = 1.11x10 ^ -5 M

pH = - log 1.11x10 ^ -5 = 4.95

b) 1.3x10 ^ -10 = X ^ 2 / 0.394

X = 7.16x10 ^ -6 M

pH = - log 7.16x10 ^ -6 = 5.15

c) Kb = [HA] * [OH] / [A-]

1.7x10 ^ -9 = X ^ 2 / 0.869

It clears X = 3.84x10 ^ -5 M

POH and pH are calculated:

pOH = - log 3.84x10 ^ -5 = 4.42

pH = 14 - 4.42 = 9.58

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.